### v2

@@ 4 Discovering Anchored Bias in MCQs
As introduced in § 3, we use 5 different datasets with choices from 2 to 5 to investigate the anchored bias. Table 1 shows that the correct choices distribution of ARC and CSQA training dataset is balanced from ‘A’ to ‘E’. In addition, IOI, LD and Greater test datasets are manually synthesised, and we did not choose or place the correct choice at ‘A’. For all test datasets, we only select samples whose correct choice is not ‘A’ to avoid introducing extra bias, i.e., the mix-up between correct prediction and anchored bias of GPT2 family. Based on the randomly sampled test datasets in Table 1, we further calculate the distribution of anchored bias ‘A’ happened within each test dataset when different GPT2 model is used in Table 2. We can find that GPT2-Large and GPT2-Medium have the most serious anchored bias, and GPT2-XL and GPT2-Small have relatively less serious issues. Based on this situation, we mainly focus on investigating test samples which have anchored bias for each dataset. Table 7 in Appendix E shows the number of test samples for each dataset and GPT2 model, where Infer. is used to localise and mitigate anchored bias and Eva. is used to verify the performance of mitigation.
@@
We first investigate MLP modules within GPT2 family for anchored bias. Inspired from Geva et al., (2021, 2022), the MLP modules can be regarded as key-value memories. As introduced in § 2 and Fig. 2, the keys of MLP module is a vector of coefficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, which dynamically controls the contributions of the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$ based on different input prompts. The value $\mathbf{v}_{i}^{\ell}$ is treated as a memory bank which stores knowledge after the model pertaining.
@@
We use logit lens (Nostalgebraist,, 2020) to calculate logit of the final input prompt token contributing to incorrect choice token ‘A’ and correct choice token ‘B/C/D/E’ based on different datasets using Eq. 4 and Eq. 5333The reason why we focus on the final input token is that the information inside of autoregressive transformer-based model will accumulate to the final input token for the next token prediction.:
$\text{logit}_{T}^{\ell}(\mathbf{m}_{T}^{\ell})[\texttt{A}]=W_{U}[\texttt{A}]\sigma(\mathbf{m}_{T}^{\ell})\quad\text{logit}_{T}^{\ell}(\mathbf{m}_{T}^{\ell})[\texttt{B/C/D/E}]=W_{U}[\texttt{B/C/D/E}]\sigma(\mathbf{m}_{T}^{\ell})$ (6)
@@
We calculate the MLP logit difference between anchored bias token ‘A’ and correct choice token B/C/D/E averaged across all layers and datasets for each GPT2 model using Infer. test samples. As shown in Fig. 3, layer 9 in GPT2-Small, layer 20 in GPT2-Medium, layer 34 in GPT2-Large, and layer 37/38/44 in GPT2-XL are dominant layers [^4] In this work, the layer number and head number start from 0. related to anchored bias. In addition, these layers are much closer to the final layer of GPT2, which agrees with Geva et al., (2022); Gurnee et al., (2023)’s finding that higher layers in GPT2 are relevant to semantic concepts or complicated tasks. We also notice that the last one or two layers in each GPT2 model do not have anchored bias at all, and they contribute more logits to correct choice token ‘B/C/D/E’ to ‘A’. However, as the anchored bias logits are accumulated from previous layers, the final one or two layers cannot totally correct this bias.
Based on the pattern from Fig. 3, we further use MLP contribution (Geva et al.,, 2022) to localise the specific dimensionality from $W_{\text{out}}^{\ell}$ in these identified layers leading to anchored bias:
@@
In the Table 3, vector column demonstrates the specific value vectors $\mathbf{v}^{\ell,n}$ which are responsible for anchored bias. For each value vector, we also calculate how frequently it is recognised as an anchored-bias vector across all datasets for each GPT2 model. We can find that most identified value vectors have more than 50% chance with anchored bias happening across all datasets and different GPT2 models. To further verify whether these value vectors store anchored knowledge bias, we unembeded each value vector logit and selected the top-10 tokens with the highest probability. As shown in Table 3, we can find that most top-10 tokens within each value vector are relevant to ‘A’, e.g., ␣A, A, ␣a, a, etc. This finding proves that some value vectors in $W_{\text{out}}$ of GPT2 family store knowledge bias after pertaining, and these knowledge biases will become anchored bias when the input prompt is formatted as MCQ. In addition, most unembeded tokens are stopwords, e.g., pronouns, articles, prepositions, etc, which also agrees with the findings that “stopwords/punctuation” are commonly distributed in the value vectors of MLP (Geva et al.,, 2022).
Table 3: Identified anchored-bias value vectors $\mathbf{v}^{\ell,n}$ of $n$-row of $W_{\text{out}}^{\ell}$ at layer $\ell$ for each GPT2 model, where the percentage indicates how frequently the specific $\mathbf{v}^{\ell,n}$ is detected as an anchored-bias vector across all datasets, and represents single space within the token because GPT2 tokeniser encodes same word with or without as different token numbers. For each value vector, we further unembedded the top-10 tokens, and most of them are human-interpretable words, which also verify that pretrained GPT2 family has intrinsic anchored bias within $W_{\text{out}}$ (See Appendix F for more unembedded tokens for each GPT2 model).
@@
Following a similar method as locating anchored bias in MLP, we also aim to solve these research questions: 1) Is the attention head also responsible for the anchored bias in GPT2 family? 2) Which layer and head of attention pattern is anchored bias relevant to?
Table 4: The new top-10 tokens of each updated value vector for each GPT2 model (See Appendix G for more new unembedded tokens for each GPT2 model).
@@ 5 Mitigating Anchored Bias in MCQs
Figure 6: The full circuit of anchored bias for GPT2-Small model, where each attention head and MLP module are selected when MLP and attention pattern logit difference threshold is larger than 4. The percentage within each module indicates the probability of anchored bias across different datasets for GPT2-Small model when threshold is larger than 4 (See Appendix M for the full circuits of other GPT2 models).
@@
Following (Dai et al.,, 2022), we directly modify and update the identified value vector as:
@@
where $\lambda_{1}=1,\lambda_{2}=8$. After updating the corresponding value vector in MLP, we utilise the updated GPT2 model to predict the next token of the same input MCQ prompts. We comprehensively evaluate each updated GPT2 model with the corresponding modified value vector using Infer. and Eva. across all datasets. As shown in Table 5, most updated value vectors achieve high classification accuracy regarding MCQ tasks, even with multiple near 100% or 100% accuracy in $\mathbf{v}^{9,1853}$ of GPT2-Small, $\mathbf{v}^{34,1541}$ of GPT2-Large, $\mathbf{v}^{44,4967}$ and $\mathbf{v}^{38,4191}$ of GPT2-XL. For those value vectors with around a 60-70% chance of anchored bias happening across all datasets and different GPT2 models (i.e., $\mathbf{v}^{9,2859}$ of GPT2-Small, $\mathbf{v}^{20,3713}$ of GPT2-Medium, $\mathbf{v}^{34,2103}$ of GPT2-Large, and $\mathbf{v}^{37,2966}$ of GPT2-XL in Appendix H), the classification accuracy still 68.09% averaged all datasets and all models. This means that the simple and straightforward method (i.e., Eq. 8) is effective, and we do not need to fine-tune the whole GPT2 model to fix the anchored bias. We further unembedded the updated value vectors in Table 4, and it shows that the anchored bias token ‘A’ is significantly removed and the new top-10 tokens for each new value vector are replaced with the correct choices token ‘B/C/D/E’.
Table 5: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=8$ (See Appendix H for the ablation study of $\lambda_{2}$ with different values).
@@
Table 6: The Top-1 classification accuracy changes after attention pattern recalibration for each GPT2 model across all datasets (See Appendix I for other GPT2 model’s results).
@@ 6 Discussion
Based on the comprehensive zero-shot learning MCQ across all datasets, we have a good understanding of how important the MLP and attention head are regarding the anchored bias in the GPT2 family. The following question might be whether few-shot learning can mitigate this anchored bias without updating specific value vectors in MLP or recalibrating the attention head. We conduct an experiment to evaluate GPT2 family across all datasets using 1-shot and 2-shot learning settings. The initial finding is that the anchored bias could be relatively mitigated and average MCQ classification accuracy across GPT2 family is 46.74%, 44.44%, 38.46% and 23.34% under 1-shot learning and 46.52%, 45.70%, 43.22% and 32.62% under 2-shot learning (See Appendix J). This result indicates that GPT2 family still struggles to predict correct choices, especially for GPT2-XL, which needs more investigation in the future.
@@
We conduct an experiment to evaluate whether direct value vector updating in MLP is harmful to the general ability of GPT2-Small for the original IOI and Greater-than tasks. The experiment shows that the modified GPT2 family can still achieve the average 85.8%, 91.1%, 81.7% and 78.8% accuracy on the original IOI dataset, and 96.1%, 98.0%, 98.5% and 98.4% on the original Greater-than dataset (See Appendix K). Although direct value vector updating in MLP is harmful to the general ability of GPT2 on the original IOI and Greater-than datasets, this model editing approach does not produce serious damage, which matches the findings from Gu et al., (2024). However, we need to develop a better and minimal-harm model editing algorithm in the future.
@@
We construct two random MCQ datasets which include random concatenated characters and random vocabulary words (see Table 18 and Table 19 in Appendix L). The result shows that anchored bias still happens across different GPT2 models, especially for GPT2-Large and GPT2-XL. This indicates that anchored bias is insensitive to MCQ input prompts, which also confirms our findings that GPT2 family exhibit the anchored bias with significant regularity across various MCQ datasets.
7 Related Work
Mechanistic Interpretability.
Our work is part of Mechanistic Interpretability research, which seeks to decode the internal mechanisms of neural networks (Bereska and Gavves,, 2024; Ferrando et al.,, 2024). Recent studies have deepened our understanding of how transformers manage factual associations. Geva et al., (2023) explores information flow during inference, while Meng et al., 2022a examines knowledge storage in GPT-like models, highlighting their distinct features and using causal mediation analysis to investigate fact recall. Further, Meng et al., 2022b develops an algorithm to enhance memory recall in transformers. Gupta et al., (2023) and Halawi et al., (2023) introduce methods for model editing and behavior analysis, respectively, to mitigate harmful behaviors and adjust commonsense knowledge. Wang et al., (2022) identifies key circuits in attention heads. Our research extends these studies by reformulating IOI as a binary question-answering task to explore positional bias in LLMs.
Positional Bias of LLMs in MCQ.
Several studies have documented the effects of positional bias on LLM accuracy in MCQs. Pezeshkpour and Hruschka, (2023) found a "sensitivity gap" in models like GPT-4, where positional bias can decrease performance by up to 75% in a zero-shot setting, and they improved accuracy with new calibration strategies. Wang et al., (2023) also noted the impact of option order on GPT-4’s scores, enhancing accuracy through a calibration framework including multiple evidence and balanced position adjustments, along with human involvement. Zheng et al., 2024a addressed "selection bias" where LLMs disproportionately favour certain options, introducing a debiasing method, PriDe, that adjusts predictions during inference. Wang et al., (2024) explored performance changes from reordering answer options, confirming that this impacts understanding. Turpin et al., (2024) and Zheng et al., 2024b explore the effects of positional bias in LLMs, showing how it skews Chain-of-Thought generation and evaluator judgments, and emphasize the need for strategies to detect and mitigate these biases. These findings highlight the need for effective strategies to mitigate positional biases, thus enhancing the reliability of LLMs in assessments.
@@ 8 Conclusion
Appendix A Limitations
This work mainly focuses on the mechanistic analysis of GPT2 family with the model size from 124M to 1.5B. It is worth comprehensively investigating whether larger open-source LLMs have similar anchored biases, such as LLaMA-7B-65B, LLaMA2-7B-70B, LLaMA3-8B-71B, etc. In addition, different LLM architectural backbones might have different extents of anchored bias, e.g., Mixture of Experts (MoE) and Mamba with selective state spaces. It is meaningful to compare how different MLPs and attention heads are across different LLMs above and why anchored bias disappears if larger LLMs do not have such an issue. Moreover, the knowledge editing approach by directly updating value vectors from MLPs is not optimal as it will introduce some extent of damage to the general ability of GPT2 models. However, how to develop a better and minimal-harm model editing algorithm is an open question (Gu et al.,, 2024), which is worth exploring in the future.
Appendix B Broader Impacts

### v3

@@ 5 Discovering Anchored Bias in MCQs
As introduced in § 4, we use 5 different datasets with choices from 2 to 5 to investigate the anchored bias. Table 1 shows that the correct choices distribution of ARC and CSQA training datasets is balanced from ‘A’ to ‘E’. In addition, IOI, LD and Greater test datasets are manually synthesised, and we did not choose or place the correct choice at ‘A’. For all test datasets, we only select samples whose correct choice is not ‘A’ to avoid introducing extra bias, i.e., the mix-up between correct prediction and anchored bias of GPT2 family. Based on the randomly sampled test datasets in Table 1, we further calculate the distribution of anchored bias ‘A’ that happened within each test dataset when different GPT2 model is used in Table 2. We can find that GPT2-Large and GPT2-Medium have the most serious anchored bias, and GPT2-XL and GPT2-Small have relatively less serious issues. Based on this situation, we mainly focus on investigating test samples which have anchored bias for each dataset. Table 8 in Appendix D shows the number of test samples for each dataset and GPT2 model, where Infer. is used to localise and mitigate anchored bias and Eva. is used to verify the performance of mitigation.
@@
We first investigate MLP modules within GPT2 family for anchored bias. Inspired from Geva et al. (2021, 2022), the MLP modules can be regarded as key-value memories. As introduced in § 3 and Fig. 2, the keys of MLP module is a vector of coefficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, which dynamically controls the contributions of the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$ based on different input prompts. The value $\mathbf{v}_{i}^{\ell}$ is treated as a memory bank which stores knowledge after the model pertaining.
@@
We use logit lens (Nostalgebraist, 2020) to calculate logit of the final input prompt token contributing to incorrect choice token ‘A’ and correct choice token ‘B/C/D/E’ based on different datasets using Eq. 4 and Eq. 5333The reason why we focus on the final input token is that the information inside of autoregressive transformer-based model will accumulate to the final input token for the next token prediction.:
$\begin{split}\text{logit}_{T}^{\ell}(\mathbf{m}_{T}^{\ell})[\texttt{A}]&=W_{U}[\texttt{A}]\sigma(\mathbf{m}_{T}^{\ell})\\ \text{logit}_{T}^{\ell}(\mathbf{m}_{T}^{\ell})[\texttt{B/C/D/E}]&=W_{U}[\texttt{B/C/D/E}]\sigma(\mathbf{m}_{T}^{\ell})\end{split}$ (6)
@@
We calculate the MLP logit difference between anchored bias token ‘A’ and correct choice token B/C/D/E averaged across all layers and datasets for each GPT2 model using Infer. test samples. As shown in Fig. 3, layer 9 in GPT2-Small, layer 20 in GPT2-Medium, layer 34 in GPT2-Large, and layer 37/38/44 in GPT2-XL are dominant layers [^4] In this work, the layer number and head number start from 0. related to anchored bias. In addition, these layers are much closer to the final layer of GPT2, which agrees with Geva et al. (2022); Gurnee et al. (2023)’s finding that higher layers in GPT2 are relevant to semantic concepts or complicated tasks. We also notice that the last one or two layers in each GPT2 model do not have anchored bias at all, and they contribute more logits to the correct choice token ‘B/C/D/E’ than to ‘A’. However, as the anchored bias logits are accumulated from previous layers, the final one or two layers cannot totally correct this bias.
Based on the pattern from Fig. 3, we further use MLP contribution (Geva et al., 2022) to localise the specific dimensionality from $W_{\text{out}}^{\ell}$ in these identified layers leading to anchored bias:
@@
Table 3: Identified anchored-bias value vectors $\mathbf{v}^{\ell,n}$ of $n$-row of $W_{\text{out}}^{\ell}$ at layer $\ell$ for each GPT2 model, where the percentage indicates how frequently the specific $\mathbf{v}^{\ell,n}$ is detected as an anchored-bias vector across all datasets, and represents single space within the token because GPT2 tokeniser encodes same word with or without as different token numbers. For each value vector, we further unembedded the top-10 tokens, and most of them are human-interpretable words, which also verify that pretrained GPT2 family has intrinsic anchored bias within $W_{\text{out}}$ (See Appendix E for more unembedded tokens for each GPT2 model).
In the Table 3, the vector column demonstrates the specific value vectors $\mathbf{v}^{\ell,n}$ which are responsible for anchored bias. For each value vector, we also calculate how frequently it is recognised as an anchored-bias vector across all datasets for each GPT2 model. We can find that most identified value vectors have more than a 50% chance with anchored bias happening across all datasets and different GPT2 models. To further verify whether these value vectors store anchored knowledge bias, we unembedded each value vector logit and selected the top-10 tokens with the highest probability. As shown in Table 3, we can find that most top-10 tokens within each value vector are relevant to ‘A’, e.g., ␣A, A, ␣a, a, etc. This finding proves that some value vectors in $W_{\text{out}}$ of GPT2 family store knowledge bias after pertaining, and these knowledge biases will become anchored bias when the input prompt is formatted as MCQ. In addition, most unembedded tokens are stopwords, e.g., pronouns, articles, prepositions, etc, which also agrees with the findings that “stopwords/punctuation” are commonly distributed in the value vectors of MLP (Geva et al., 2022).
@@
Following a similar method to locating anchored bias in MLP, we also aim to solve these research questions: 1) Is the attention head also responsible for the anchored bias in GPT2 family? 2) Which layer and head of attention pattern is anchored bias relevant to?
@@
Table 4: The new top-10 tokens of each updated value vector for each GPT2 model (See Appendix F for more new unembedded tokens for each GPT2 model).
@@ 6 Mitigating Anchored Bias in MCQs
Figure 6: The full circuit of anchored bias for GPT2-Small model, where each attention head and MLP module are selected when MLP and attention pattern logit difference threshold is larger than 4. The percentage within each module indicates the probability of anchored bias across different datasets for GPT2-Small model when threshold is larger than 4 (See Appendix L for the full circuits of other GPT2 models).
@@
[…]
where $\lambda_{1}=1,\lambda_{2}=8$. After updating the corresponding value vector in MLP, we utilise the updated GPT2 model to predict the next token of the same input MCQ prompts. We comprehensively evaluate each updated GPT2 model with the corresponding modified value vector using Infer. and Eva. across all datasets. As shown in Table 5, most updated value vectors achieve high classification accuracy regarding MCQ tasks, even with multiple near 100% or 100% accuracy in $\mathbf{v}^{9,1853}$ of GPT2-Small, $\mathbf{v}^{34,1541}$ of GPT2-Large, $\mathbf{v}^{44,4967}$ and $\mathbf{v}^{38,4191}$ of GPT2-XL. For those value vectors with around a 60-70% chance of anchored bias happening across all datasets and different GPT2 models (i.e., $\mathbf{v}^{9,2859}$ of GPT2-Small, $\mathbf{v}^{20,3713}$ of GPT2-Medium, $\mathbf{v}^{34,2103}$ of GPT2-Large, and $\mathbf{v}^{37,2966}$ of GPT2-XL in Appendix G), the classification accuracy still 68.09% averaged all datasets and all models. This means that the simple and straightforward method (i.e., Eq. 8) is effective, and we do not need to fine-tune the whole GPT2 model to fix the anchored bias. In addition, such performance gains can be generalised to the Eva. dataset without accessing gold labels. We further unembedded the updated value vectors in Table 4, and it shows that the anchored bias token ‘A’ is significantly removed and the new top-10 tokens for each new value vector are replaced with the correct choices token ‘B/C/D/E’.
@@
Table 5: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=8$ (See Appendix G for the ablation study of $\lambda_{2}$ with different values).
@@
Table 6: The Top-1 classification accuracy changes after attention pattern recalibration for each GPT2 model across all datasets (See Appendix H for other GPT2 model’s results).
@@ 7 Discussion
Based on the comprehensive zero-shot learning MCQ across all datasets, we have a good understanding of how important the MLP and attention head are regarding the anchored bias in the GPT2 family. The following question might be whether few-shot learning can mitigate this anchored bias without updating specific value vectors in MLP or recalibrating the attention head. We conduct an experiment to evaluate GPT2 family across all datasets using 1-shot and 2-shot learning settings. The initial finding is that the anchored bias could be relatively mitigated and average MCQ classification accuracy across GPT2 family is 46.74%, 44.44%, 38.46% and 23.34% under 1-shot learning and 46.52%, 45.70%, 43.22% and 32.62% under 2-shot learning (See Appendix I). This result indicates that GPT2 family still struggles to predict correct choices, especially for GPT2-XL, which needs more investigation in the future.
Models CNN (Rouge-Avg) WikiText2 (PPL) PTB (PPL)
GPT2-Small 16.8 / 15.2 31.2 / 33.5 66.1 / 67.4
GPT2-Medium 20.4 / 18.6 23.8 / 25.9 48.2 / 49.7
GPT2-Large 21.5 / 19.9 20.5 / 22.3 41.8 / 43.2
GPT2-XL 21.7 / 20.2 19.1 / 20.1 36.2 / 37.9
Table 7: The performance of different GPT2 models on three different tasks with random-selected 1000 samples before and after updating the identified MLP value vectors in the § 6. We use / to separate the performance before and after the intervention.
@@
We conduct an experiment to evaluate whether direct value vector updating in MLP is harmful to the general ability of GPT2-Small for the original IOI and Greater-than tasks. The experiment shows that the modified GPT2 family can still achieve the average 85.8%, 91.1%, 81.7% and 78.8% accuracy on the original IOI dataset, and 96.1%, 98.0%, 98.5% and 98.4% on the original Greater-than dataset (See Appendix J). Although direct value vector updating in MLP is harmful to the general ability of GPT2 on the original IOI and Greater-than datasets, this model editing approach does not produce serious damage, which matches the findings from Gu et al. (2024). Similarly, we further evaluate the general ability of GPT-2 families on the CNN and Daily Mail, WikiText2 and PTB tasks using the identified value vector of MLP in the § 6. Table 7 demonstrates similar findings to the Appendix J. However, we need to develop a better and minimal-harm model editing algorithm in the future.
@@
We construct two random MCQ datasets which include random concatenated characters and random vocabulary words (see Table 19 and Table 20 in Appendix K). The result shows that anchored bias still happens across different GPT2 models, especially for GPT2-Large and GPT2-XL. This indicates that anchored bias is insensitive to MCQ input prompts, which also confirms our findings that GPT2 family exhibit the anchored bias with significant regularity across various MCQ datasets.
@@ 8 Conclusion
Acknowledgement
We would like to thank all the anonymous reviewers for their insightful comments. This work is supported by the Gemma 2 Academic Program GCP Credit Award from Google.
Limitations
This work mainly focuses on the mechanistic analysis of GPT2 family with the model size from 124M to 1.5B. It is worth comprehensively investigating whether larger open-source LLMs have similar anchored biases, such as LLaMA-7B-65B, LLaMA2-7B-70B, LLaMA3-8B-71B, etc. In addition, different LLM architectural backbones might have different extents of anchored bias, e.g., Mixture of Experts (MoE) and Mamba with selective state spaces. It is meaningful to compare how different MLPs and attention heads are across different LLMs above and why anchored bias disappears if larger LLMs do not have such an issue. Moreover, the knowledge editing approach by directly updating value vectors from MLPs is not optimal as it will introduce some extent of damage to the general ability of GPT2 models. However, how to develop a better and minimal-harm model editing algorithm is an open question (Gu et al., 2024), which is worth exploring in the future.
Appendix A Broader Impacts